The number of atoms in 22g of strontium is z. What is the number of bromine molecules in 20g of bromine gas ?
1
Expert's answer
2015-11-03T12:19:29-0500
MWSr = 87.62 g/mol
MWBr2 = 159.808 g/mol

n = N/NA = m/MW;
N = m NA/MW
N(Sr)/N(Br2) = (m(Sr) NA/MW(Sr))/(m(Br2) NA/MW(Br2)) = m(Sr) MW(Br2)/m(Br2) MW(Sr) = (22 g 159.808 g/mol)/(20 g 87.62 g/mol) = 2.0

N(Br2) = 0.5 N(Sr) = 0.5 z